C5H12 (l) + 8O2 (g) → 5CO2 (g) + 6 H2O (g)


1. Given 4.3 x 1024 atoms of oxygen (O), how many moles of carbon dioxide (CO2) are present? Answer with 2 sig figs.



1
Expert's answer
2022-05-05T17:25:04-0400

By the balanced reaction the proportion of the "O_2" and "CO_2" mol number is:

"8n(O_2) = 5n(CO_2)"


By the given details, the mol number of the "O_2" can be deduced from the following formula:

"n(O_2) = \frac{N(O_2)}{N_A}" , where the "N_A= 6.022 \times10^{23} p/mol" Avogadro's number.


"n(O_2) = \frac{4.3 \times 10^{24}}{6.022 \times 10^{22}} = 0.71\times10^2 mol"


Using the first proportion from the balanced equation the mol number of the carbon dioxide is:

"n(CO_2)=\frac{8n(O_2)}{5} = \frac{8\times 0.71\times 10^2}{5} = 1.1 \times 10^2 mol"